# Fish Tracking, Counting, and Behaviour Analysis in Digital Aquaculture: A Comprehensive Review

### v1

@@ Abstract
Digital aquaculture leverages advanced technologies and data-driven methods, providing substantial benefits over traditional aquaculture practices. Fish tracking, counting, and behaviour analysis are crucial components of digital aquaculture, which are essential for optimizing production efficiency, enhancing fish welfare, and improving resource management. Previous reviews have focused on single modalities, limiting their ability to address the diverse challenges encountered in these tasks comprehensively. This review provides a comprehensive analysis of the current state of aquaculture digital technologies, including vision-based, acoustic-based, and biosensor-based methods. We examine the advantages, limitations, and applications of these methods, highlighting recent advancements and identifying critical research gaps. The scarcity of comprehensive fish datasets and the lack of unified evaluation standards, which make it difficult to compare the performance of different technologies, are identified as major obstacles hindering progress in this field. To overcome current limitations and improve the accuracy, robustness, and efficiency of fish monitoring systems, we explore the potential of emerging technologies such as multimodal data fusion and deep learning. Additionally, we contribute to the field by providing a summary of existing datasets available for fish tracking, counting, and behaviour analysis. Future research directions are outlined, emphasizing the need for comprehensive datasets and evaluation standards to facilitate meaningful comparisons between technologies and promote their practical implementation in real-world aquaculture settings.
@@
Digital aquaculture, fish tracking, counting, behaviour analysis
@@ I Introduction
Fish tracking, counting, and behaviour analysis are crucial components of digital aquaculture, playing a vital role in effective management and decision-making. Accurate monitoring of these aspects can help detect abnormal fish behaviour, estimate fish abundance, and formulate reasonable management strategies, ultimately improving fish welfare and economic outcomes in the aquaculture industry [4]. Traditional methods for fish tracking and behaviour analysis rely on the experience of human observers, and the observation results depend on their skills and knowledge, which are not always reliable [5, 6]. Similarly, manual fish counting methods involve removing fish from tanks, leading to stress, injury, and disease, negatively impacting fish welfare and growth [7, 8]. The implementation of intelligent tracking, counting, and behaviour analysis technologies can help overcome these limitations, reducing the risk of fish mortality, improving feeding strategies, and promoting sustainable development in aquaculture [9, 10, 11].
@@
Currently, various technologies such as vision-based sensors, acoustic-based sensors and biosensors methods are used for fish tracking, counting, and behaviour analysis in aquaculture. Vision-based sensors and computer vision technology have found widespread application due to advancements in optical imaging and computer vision. However, they are limited by poor illumination, low contrast, high noise, fish deformation, frequent occlusion, and dynamic backgrounds [12, 13, 14, 15]. Acoustic-based sensors and hydroacoustic methods, which are non-invasive, are particularly useful for monitoring fish in turbid water environments and overnight, but their high hardware cost limits their popularity in intensive aquaculture settings [16, 17, 18, 19, 20]. Biosensors can provide valuable information on fish physiology and behaviour, but their invasive nature and the need for individual fish tagging can be challenging in large-scale aquaculture operations [21].
Previous reviews have been conducted on fish tracking, counting and behaviour analysis [22, 23, 10, 24, 12, 25]. However, most reviews focus on computer vision technology as the primary approach and relying on a single modality may not provide sufficient data for comprehensive analysis. To address this limitation, our paper systematically surveys vision-based sensors, acoustic-based sensors, biosensors, and hydroacoustic methods, facilitating a holistic discussion of tracking, counting, and behaviour analysis while identifying technology gaps in the current literature.
@@
This article comprehensively reviews the literature on fish tracking, counting, and behavioural analysis in aquaculture over the past two decades, emphasizing the progress made in these areas and identifying potential future research directions. The remainder of this article is structured as follows: Section II explores the advancements in fish tracking techniques, while Section III discusses the various methods and applications of fish counting. Section IV discuss the behaviour analysis of fish, and Section V presents an overview of relevant public datasets. In Section VI, we examine the challenges faced by the aquaculture industry and discuss future development trends. Finally, Section VII summarises the key findings and conclusions presented in this paper.
@@ II Fish tracking
Vision-based multi-target tracking methods are increasingly used in fish behaviour analysis. However, fish tracking is challenging because of the small differences between individuals, complex environments, and variations in plankton, shapes, angles, and scales of swimming fish [26]. Fish tracking can be categorized into two-dimensional (2D) and three-dimensional (3D) tracking based on the swimming environment [27]. 2D tracking is used in shallow water containers, where fish swimming appears close to a 2D planar motion and is represented using (x, y) coordinates, but can only analyze a part of the fish behaviour. In contrast, 3D tracking considers depth information and is represented using (x, y, z) coordinates, enabling the analysis of spatial movement in natural environments.
In addition to vision-based tracking, acoustic techniques such as the Acoustic Tag System (ATS) are also used for fish tracking. ATS involves attaching acoustic tags to fish, which emit unique acoustic signals that are detected by hydrophone receivers. The position of the tagged fish can be estimated using the time difference of arrival of the acoustic signals at multiple receivers, allowing for 3D tracking of fish movement in natural habitats [28] This section mainly analyzes the relevant literature on fish tracking methods based on visual technology (as shown in Table I) and acoustic techniques in recent years and provides a systematic summary.
@@ Methods
Tank 11
Fish
center
points
YOLOV3
Euclidean
Distance
Enhance target
detection in
unclear water
Fish numbers are
too small and
done in the
laboratory
[30]
@@
Tank - Centroids Otsu
Manhattan
Distance
Low cost and
removable
installation
Stationary fish
mistaken for
debris or dead
[14]
@@
25 Fish head DOH CNN Recall
@@ II-A Fish tracking based on 2-dimensional visual information
Fish tracking methods can be broadly categorized into three main approaches: classical algorithms, kernel correlation Filter algorithms, and deep learning-based tracking algorithms [25]. Each category encompasses various techniques with their strengths and limitations, which will be explored in more detail in the following subsections.
II-A1 Fish tracking based on classical algorithm
@@
Classical algorithms have been widely used to address the challenges of fish tracking in complex underwater environments, such as rapid posture changes, occlusion, overlap, and poor image quality. The Tracking-Learning-Detection (TLD) algorithm, which updates salient features and target model parameters through online learning, has shown promise in providing stable tracking [42]. However, its median-flow tracker may fail when fish change their swimming posture rapidly. An adaptive scale mean-shift (ASMS) algorithm, utilizing fish shape and colour features, has replaced the median-flow tracker and handled posture changes, uneven illumination, and complex backgrounds [43].
@@
Preserving individual fish identities during occlusion and overlap remains a significant challenge. Techniques that extract head shape or body geometry features have been explored [44, 45], but their effectiveness may be limited by the rapid movement and intense geometry of fish bodies [14]. Adaptive threshold algorithms, which estimate thresholds for each pixel based on its adjacent region, have shown promise in segmenting individual fish in binarized images [46]. The global nearest neighbour algorithm with fish posture as a tracking constraint has been used to track small numbers of zebrafish [36], but it lacks individual recognition ability, leading to track exchanges during overlap or occlusion. The Toxld algorithm addresses this issue using intensity histograms and Hu-moments to link trajectory fragments and preserve individual fish identities [47]. However, the error increases with the number of fish.
@@
To mitigate poor image quality, retinex (MSR) based enhancement algorithms combined with object detection and Euclidean tracking have been used to improve fish detection in unclear underwater images [30, 48, 49]. Kalman Filters may not always be optimal for underwater fish tracking due to non-Gaussian noise and complex environments [50]. Mean offset technology, which models fish probability density based on colour histograms, can fail when the background colour closely resembles the fish colour distribution. Tracking algorithms based on covariance representation, which model objects as covariance matrices of pixel-based feature sets, incorporate spatial and statistical characteristics, making them more suitable for tracking fish in challenging underwater scenarios [51, 52].
II-A2 Fish tracking based on Kalman Filter
The Kalman Filter, an efficient autoregressive filter that estimates the state of a dynamic system in an environment with uncertainties, has been widely used for fish tracking due to its versatility and robustness [53, 33]. Building upon the Kalman Filter, the SORT (Simple Online and Realtime Tracking) algorithm has emerged as a simple yet effective multi-target tracking approach [34]. SORT utilizes a Kalman Filter for frame-by-frame data correlation, and the Hungarian algorithm for correlation measurement [54]. Despite its good performance at high frame rates, SORT has limitations, such as ignoring object surface features, which makes the tracking results heavily dependent on the detection performance [55]. To address this issue, an extension of SORT called DeepSORT was developed, which leverages a CNN model trained on large-scale pedestrian datasets to extract features that enhance the network’s robustness to loss and obstacle [56].
Recent literature shows that DeepSORT, an extension of the SORT algorithm, has been extensively applied in fish tracking [29]. DeepSORT combines the Kalman Filter-based SORT framework with a deep learning-based appearance feature extractor, enabling more robust tracking performance. However, challenges arise when fish undergo rapid body shape changes during fast turns, leading to blurry and difficult-to-track images [32]. To mitigate this issue, shorter exposure times and variable-size boundary boxes can be used, with the boundary boxes being estimated according to the motion state. Despite these challenges, the Kalman Filter remains a popular choice for fish tracking due to its ability to estimate the state of a dynamic system in the presence of uncertainties.
The frame rate plays a crucial role in Kalman Filter-based fish tracking performance. Low frame rates can lead to increased tracks and a higher likelihood of misidentification. Conversely, high frame rates result in more linear fish motion, enabling Kalman Filters to predict individual motion more accurately (as shown in Fig. 1). As the field continues to advance, future research should prioritize optimizing tracking schemes to minimize computing time and evaluating the long-term tracking performance of these methods in diverse aquatic environments.
Figure 1: Fish trajectory under different frame rates.
@@ II-A3 Fish tracking based on deep learning
Deep learning has emerged as a powerful tool for fish tracking, with Tracking by Detection (TBD) being the primary approach. In TBD, a deep learning model is trained on a large dataset to learn convolutional features with strong expressive power, enabling the detection and tracking of fish in video sequences. However, occlusion between fish remains a significant challenge in TBD methods, often leading to the generation of fragmented trajectories that require post-processing to link them together [57].
To address the occlusion issue, several notable multi-target tracking algorithms have been proposed, such as idTracker [58] and its upgraded version, idtracker.ai [59]. These algorithms extract unique fingerprint features from each animal in a set of videos and then identify each target in the video, enabling the tracking of individuals within a group by automatically identifying untagged animals. Although these methods have been widely used for tracking juvenile fish and small animals, the experimental setup restricted fish from swimming up and down to avoid overlapping, simplifying the task compared to real-world 3D tracking scenarios.
Further advancements in deep learning-based fish tracking have been achieved by combining CNN-based methods with other techniques, such as head detection, motion state prediction, and verification using SVM classifiers [41]. These approaches have demonstrated more robust tracking performance compared to idTracker when the fish density is higher, and the occlusion frequency increases, highlighting the potential of deep learning in handling complex tracking scenarios [35].
@@
Despite the progress made in controlled laboratory environments, real-world marine environments pose additional challenges for fish tracking, such as light fluctuations and waves. To address these issues, To tackle these issues, researchers have developed methods like the real-time multi-class fish stock statistics method (RMCF), which uses YOLOv4 as the backbone network and adopts a parallel two-branch structure based on deep learning for detecting fish species, tracking, and counting fish [31]. Although these methods have shown promising results in complex marine environments, their recognition accuracy may vary in different sea areas due to differences in colour cast and contrast, necessitating the retraining of the network weight coefficients.
@@
Siamese network trackers have gained attention in recent years due to their exceptional tracking speed and high accuracy. The introduction of advanced algorithms, such as SiamRPN++ (as shown in Fig. 2), has further highlighted the performance of Siamese networks, surpassing the performance of tracking algorithms based on correlation filters [60], [61]. Although there are currently few articles on Siamese networks specifically for fish tracking, this approach is expected to become a new direction in the field.
@@
Figure 2: Fish tracking method based on YoloV5 and SiamRPN++ [61].
Moreover, the emergence of transformer-based tracking methods has revolutionized the field of object tracking. Initially proposed for natural language processing tasks, transformers have been successfully adapted for computer vision tasks, including object detection and tracking [62]. Transformer-based trackers, such as TransTrack [63] and STARK [64], have demonstrated state-of-the-art performance on various tracking benchmarks. Transformer-based tracking methods have also shown promising results in fish tracking applications [39]. As transformer-based methods continue to advance in object tracking, they are expected to play an increasingly important role in fish-tracking applications. Future research should focus on further adapting transformer architectures to the specific challenges of underwater environments and developing efficient training strategies to handle the limited availability of annotated fish-tracking datasets.

### v2

# Fish Tracking, Counting, and Behaviour Analysis in Digital Aquaculture: A Comprehensive Survey

@@ Abstract
Digital aquaculture leverages advanced technologies and data-driven methods, providing substantial benefits over traditional aquaculture practices. This paper presents a comprehensive review of three interconnected digital aquaculture tasks, namely, fish tracking, counting, and behaviour analysis, using a novel and unified approach. Unlike previous reviews which focused on single modalities or individual tasks, we analyse vision-based (i.e. image- and video-based), acoustic-based, and biosensor-based methods across all three tasks. We examine their advantages, limitations, and applications, highlighting recent advancements and identifying critical cross-cutting research gaps. The review also includes emerging ideas such as applying multi-task learning and large language models to address various aspects of fish monitoring, an approach not previously explored in aquaculture literature. We identify the major obstacles hindering research progress in this field, including the scarcity of comprehensive fish datasets and the lack of unified evaluation standards. To overcome the current limitations, we explore the potential of using emerging technologies such as multimodal data fusion and deep learning to improve the accuracy, robustness, and efficiency of integrated fish monitoring systems. In addition, we provide a summary of existing datasets available for fish tracking, counting, and behaviour analysis. This holistic perspective offers a roadmap for future research, emphasizing the need for comprehensive datasets and evaluation standards to facilitate meaningful comparisons between technologies and to promote their practical implementations in real-world settings.
@@
Digital aquaculture, fish tracking, counting, behaviour analysis, multimodal fusion
@@ I Introduction
As enabling technologies in digital aquaculture, fish tracking, counting, and behaviour analysis have been studied extensively. Traditional methods for fish tracking and behaviour analysis rely on the experience of human observers, and the observation results depend on their skills and knowledge, which are not always reliable [4, 5]. Similarly, manual fish counting methods involve removing fish from tanks, leading to stress, injury, and disease, negatively impacting fish welfare and growth [6, 7]. The implementation of intelligent tracking, counting, and behaviour analysis technologies can help overcome these limitations, reducing the risk of fish mortality, improving feeding strategies, and promoting sustainable development in aquaculture [8, 9, 10]. This paper aims to provide a comprehensive survey of the research problems, existing technologies, their connections, and potential gaps in this field.
There are already literature reviews conducted previously in this field, such as [11, 12, 9, 13, 14, 15]. However, most of them focused on individual tasks or single modalities, primarily emphasizing computer vision technology as the main approach. In practice, these tasks are interconnected and form a coherent part of an overall digital aquaculture system. For example, accurate monitoring of these tasks is vital for detecting abnormal fish behaviour, estimating fish abundance, and formulating effective management strategies, ultimately improving fish welfare and economic outcomes [16]. The narrow focus in previous reviews is inherently limited in capturing the interconnected nature of these tasks in real-world aquaculture scenarios. Different from these reviews, our focus is not only on surveying the progress made for these tasks, but also on establishing the connections among them.
@@
Various technologies such as vision-based, acoustic-based and biosensing-based methods are used for fish tracking, counting, and behaviour analysis in aquaculture. Vision-based sensors and computer vision technology have found widespread application due to advancements in optical imaging and computer vision. However, they are limited by poor illumination, low contrast, high noise, fish deformation, frequent occlusion, and dynamic backgrounds [14, 17, 18, 19]. Acoustic-based sensors and hydroacoustic methods, which are non-invasive, are beneficial for monitoring fish in turbid water environments and overnight, but their high hardware cost limits their popularity in intensive aquaculture settings [20, 21, 22, 23, 24]. Biosensors can provide valuable information on fish physiology and behaviour, but their invasive nature and the need for individual fish tagging can be challenging in large-scale aquaculture operations and may potentially harm the fish [25].
A systematic review of literature from 2000 to 2023 reveals significant growth in research related to these technologies. Annual publications have increased from around 200 in the early 2000s to over 2,000 in recent years. Computer vision techniques dominate the field, accounting for about 40% of the publications, followed by acoustic monitoring methods and behavioural analysis approaches, each representing about a quarter of the research outputs. The geographical distribution of research and methodological preferences show interesting patterns. Asia, particularly China and Japan, leads in publications, contributing nearly half of the global research outputs, mostly on computer vision techniques in tracking and counting, driven by the prevalence of large-scale intensive aquaculture operations and robust technological infrastructure. Europe, contributing about 30% of publications, emerges as a leader in acoustic monitoring and behavioural analysis techniques. This trend aligns with the region’s stringent welfare regulations and emphasis on environmental sustainability in aquaculture practices. North America, while producing fewer publications, pioneers the integration of machine learning with various monitoring methods, leveraging its strong artificial intelligence (AI) research base and focus on automation. However, this global research landscape also reveals significant gaps. Developing regions, including parts of Africa, Asia, and Latin America, show limited adoption of advanced monitoring technologies. This gap can be attributed to the high initial costs, the lack of technical expertise, and the predominance of small-scale farming in these areas. Oceania, despite its significant aquaculture industry, is underrepresented in research outputs, possibly due to a smaller research community and a focus on practical applications over academic publications. Eastern Europe and Central Asia also show fewer publications on novel monitoring technologies, likely due to the later development of intensive aquaculture in these regions and limited research funding.
To address these limitations, our paper takes a holistic approach by systematically surveying a range of technologies, including vision-based sensors, acoustic-based sensors, biosensors, and hydroacoustic methods. We consider fish tracking, counting, and behaviour analysis as interconnected components of a unified task in digital aquaculture. This comprehensive perspective facilitates a more integrated discussion of these tasks while identifying technology gaps in the current literature. Furthermore, we explore the potential of emerging technologies such as multimodal data fusion, deep learning, and multi-task learning models that can simultaneously handle multiple aspects of fish monitoring. We also discuss the integration of large language models (LLM) and their potential to enhance understanding and decision-making in digital aquaculture. By adopting this multifaceted approach, our review aims to provide a more complete understanding of the challenges and opportunities in digital aquaculture, paving the way for developing more efficient, accurate, and integrated fish monitoring systems. The insights provided in this review have the potential to guide future research directions, inform policy decisions, and ultimately contribute to the sustainable growth of the aquaculture industry.
@@
The remainder of this article is structured as follows: Section II explores the advancements in fish tracking techniques, while Section III discusses the various methods and applications of fish counting. Section IV discusses the behaviour analysis of fish, and Section V introduces the potential of multimodal fusion in aquaculture. Section VI presents an overview of relevant public datasets. In Section VII, we examine the challenges faced by the aquaculture industry and discuss future development trends. Finally, Section VIII summarises the key findings and conclusions presented in this paper.
@@ II Fish tracking
Vision-based multi-target tracking methods are increasingly used in fish behaviour analysis. However, fish tracking is challenging because of the small differences between individuals, complex environments, and variations in plankton, shapes, angles, and scales of swimming fish [26]. Fish tracking can be categorized into two-dimensional (2D) and three-dimensional (3D) tracking based on the swimming environment [27]. 2D tracking is used in shallow water containers, where fish swimming appears close to a 2D planar motion and is represented using (x, y) coordinates. Still, it can only analyse a part of the fish’s behaviour. In contrast, 3D tracking considers depth information and is represented using (x, y, z) coordinates, enabling the analysis of spatial movement in natural environments.
In addition to vision-based tracking, acoustic techniques such as the Acoustic Tag System (ATS) are also used for fish tracking. ATS involves attaching acoustic tags to fish, emitting unique acoustic signals detected by hydrophone receivers. The position of the tagged fish can be estimated using the time difference of arrival of the acoustic signals at multiple receivers, allowing for 3D tracking of fish movement in natural habitats [28]. This section mainly analyzes the relevant literature on fish tracking methods based on visual technology (as shown in Table I) and acoustic techniques in recent years and provides a systematic summary.
@@ Methods
25 Fish head DOH3 CNN Recall
@@
  • 1
    CTR (correct tracking rate), CIR (correct identification reason)
  • 2
    MOTA (multiple objects tracking accuracy) (as shown in formula 1)
  • 3
    DOH (determinant of Hessian)
  • 4
    GNN (global nearest neighbour)
  • 5
    IDF1(identification-score, as shown in formula 3)
@@ II-A Fish tracking based on 2-dimensional visual information
Fish tracking methods can be broadly categorized into three main approaches: classical algorithms, kernel correlation filter-based algorithms, and deep learning-based tracking algorithms [15]. Each category encompasses various techniques with their strengths and limitations, which will be explored in more detail in the following subsections.
II-A1 Fish tracking based on classical algorithms
@@
Classical algorithms have been widely used to address the challenges of fish tracking in complex underwater environments, such as rapid posture changes, occlusion, overlap, and poor image quality. The Tracking-Learning-Detection (TLD) algorithm, which updates salient features and target model parameters through online learning, has shown promise in providing stable tracking [41]. However, its median-flow tracker may fail when fish change their swimming posture rapidly. However, an adaptive scale mean-shift (ASMS) algorithm, utilizing fish shape and colour features, can handle posture changes, uneven illumination, and complex backgrounds [42].
@@
Preserving individual fish identities during occlusion and overlap remains a significant challenge. Techniques that extract head shape or body geometry features have been explored [43, 44], but their effectiveness may be limited by the rapid movement and intense geometry of fish bodies [18]. Adaptive thresholding algorithms, which estimate each pixel’s threshold based on its adjacent region, have shown promise in segmenting individual fish in binarized images [45]. The global nearest neighbour algorithm with fish posture as a tracking constraint has been used to track small numbers of zebrafish [35], but it lacks individual recognition ability, leading to track exchanges during overlap or occlusion. The Toxld algorithm addresses this issue using intensity histograms and Hu-moments to link trajectory fragments and preserve individual fish identities [46]. However, with this method, the error increases with the number of fish.
@@
To deal with poor image quality, retinex (MSR) based enhancement algorithms combined with object detection have been used to improve fish detection in unclear underwater images [47, 48, 49]. Kalman filters may not always be optimal for underwater fish tracking due to the presence of non-Gaussian noise and complex environments [50]. Mean offset technology, which models fish probability density based on colour histograms, can fail when the background colour closely resembles the fish colour distribution. Tracking algorithms based on covariance matrices of pixel-based feature sets, incorporate spatial and statistical characteristics, making them more suitable for tracking fish in challenging underwater environments [51, 52].
II-A2 Fish tracking based on Kalman filters
The Kalman filter is an efficient autoregressive filter, which can estimate the state of a dynamic system from sensory measurements taken from an environment with uncertainties posed by noise and interferences [53]. However, when applied alone to complex tracking scenarios, it faces challenges in handling occlusions and maintaining accurate tracking in high-density situations [32]. To address these limitations, researchers explored combining Kalman filtering with advanced object detection methods. You Only Look Once (YOLO) [54] has emerged as a leading detection method due to its real-time performance and accuracy in identifying fish within video frames. Building upon these foundations, the Simple Online and Realtime Tracking (SORT) algorithm [33] integrates YOLO-based detection results with Kalman filtering for motion prediction and employs the Hungarian algorithm for data association [55]. This integration enables high-speed tracking (260 FPS) by focusing solely on bounding box information, making it ideal for applications requiring fast and efficient tracking. However, the performance of SORT is limited by its reliance on pure detection results without considering object appearance features, making it vulnerable to tracking failures during occlusions or complex scenarios [56].
To address these limitations, DeepSORT was developed by incorporating a convolutional neural network (CNN) model trained on large-scale datasets [57]. This advancement significantly improved tracking robustness to target loss and occlusions, making it particularly suitable for aquaculture applications [29]. For instance, Mathias et al. [54] successfully demonstrated this capability by implementing a hybrid adaptive DeepSORT with YOLOv3 [58] for underwater object tracking. However, challenges arise when fish undergo rapid body shape changes during fast turns, leading to blurry and difficult-to-track images [31]. To mitigate this issue, shorter exposure times and boundary boxes of variable size can be used, with the boundary boxes being estimated according to the motion state. Additionally, frame rate optimization has proven crucial for Kalman filter-based tracking performance: higher rates facilitate more linear fish motion predictions, while lower rates increase misidentification risks (as shown in Fig. 1).
Recent advancements in tracking algorithms have introduced several innovative approaches to address challenges in occlusions, rapid movements, and varying environmental conditions. One major category focuses on feature enhancement strategies: StrongSORT improved upon DeepSORT by incorporating stronger appearance features through a more sophisticated feature extractor based on ResNet-50 backbone networks and enhanced motion cues with Kalman filter refinements [59]. Its integration with GN-YOLOv5 [60] further optimized both performance and processing speed through better feature extraction and more accurate object detection [61]. Another category advances association strategies: BoT-SORT innovated tracking through a redesigned association mechanism that combines cosine distance metrics (comparing new detections’ features with stored tracklet features) for appearance-matching and Camera Motion Compensation (CMC) for handling camera movements [62]. Unlike StrongSORT which primarily relies on appearance features, BoT-SORT introduced a hybrid association strategy that balances appearance similarity with motion consistency. This dual approach has proven particularly effective in challenging scenarios like aquaculture, where combined with YOLOv8 [63], it successfully tracks fast-moving fish despite rapid movements and appearance changes by better leveraging both motion and appearance information [64].
The third category introduces observation-centric (OC) approaches, where OC-SORT and Deep OC-SORT target dense population scenarios by prioritizing actual detection results over predictions, unlike traditional trackers that rely heavily on predictive models [65, 66]. OC-SORT prioritizes high-quality observations and introduces an observation-centric Kalman filter (OCKF) for more reliable state estimation, along with a new track recovery strategy and adaptive track confidence mechanism [65]. Deep OC-SORT further enhances this approach by incorporating a deep association module for better feature matching, an enhanced track rebirth mechanism, and more sophisticated appearance modelling [66]. Building on these approaches, ByteTrack pioneered a detection utilization strategy that processes both high and low-confidence detections through separate association pipelines [67]. This differs from previous approaches that typically discarded low-confidence detections, enabling ByteTrack to maintain tracking continuity even under challenging conditions. Its success in aquaculture applications [68, 69] demonstrates robust performance under varying lighting conditions while effectively reducing object loss and trajectory fragmentation, with its flexible architecture supporting various model sizes for different application scenarios [70].
The evolution of tracking algorithms from SORT to recent approaches like StrongSORT, BoT-SORT, OC-SORT, and ByteTrack offers diverse solutions for different tracking scenarios. For practical applications, StrongSORT is suitable when strong feature extraction is crucial, BoT-SORT works well with moving cameras and appearance changes, while OC-SORT and ByteTrack are particularly effective in crowded scenes. Future research could focus on combining these strengths, such as integrating robust feature extraction with observation-centric approaches, while also addressing computational efficiency to enable real-time tracking in resource-constrained environments.
@@ II-A3 Fish tracking based on deep learning
Deep learning has revolutionized fish tracking approaches, introducing various methodologies to address the complex challenges in aquaculture environments. These approaches can be broadly categorized into three main streams: (1) identity-based methods like idTracker [71] that focus on extracting unique fingerprint features for individual fish recognition, (2) detection-tracking frameworks that combine CNN-based detection with sophisticated tracking mechanisms, and (3) advanced tracking architectures, such as Siamese networks that focus on similarity learning between frames and transformers that leverage self-attention mechanisms. While these methods have shown significant improvements over traditional approaches, they each face distinct challenges in handling occlusions, maintaining tracking stability in dense populations, and adapting to varying environmental conditions [41, 72].
Notable identity-based tracking algorithms such as idTracker [71] and its upgraded version, idtracker.ai [73], have demonstrated success in individual fish tracking. idTracker pioneers the concept of extracting unique visual fingerprints for each animal, utilizing image-based features to maintain individual identities throughout videos. The upgraded version, idTracker.ai, enhances this approach by incorporating deep learning techniques. These algorithms extract unique fingerprint features from each animal and identify targets throughout the video, enabling automated tracking of untagged animals within groups. However, their application is often limited to controlled environments where fish movement is restricted to prevent overlapping, making them less suitable for real-world 3D tracking scenarios.
@@
Detection-based tracking frameworks have achieved significant advancements by combining CNN-based detection methods with complementary tracking techniques. These approaches integrate with other techniques, such as head detection, motion state prediction, and SVM-based verification [40], showing improved robustness compared to idTracker in scenarios with higher fish density and increased occlusion frequency [34]. Despite the progress made in controlled laboratory environments, real-world marine environments pose additional challenges for fish tracking, such as light fluctuations and waves. To address these issues, researchers have developed methods like the real-time multi-class fish (RMCF) stock statistics method, which uses YOLOv4 [74] as the backbone network and adopts a parallel two-branch structure based on deep learning for detecting fish species, tracking, and counting fish [30]. Although these methods have shown promising results in complex marine environments, their recognition accuracy may vary in different sea areas due to differences in colour cast and contrast, necessitating the retraining of the network weight coefficients.
@@
Advanced tracking architectures in fish tracking have evolved along two main directions: Siamese networks and transformer-based methods [75, 76]. Siamese network trackers have gained attention in recent years due to their exceptional tracking speed and high accuracy. The introduction of advanced algorithms, such as SiamRPN++ (as shown in Fig. 2), has further demonstrated the performance of Siamese networks, surpassing the performance of the tracking algorithms based on correlation filters [77], [78]. Although there are currently few articles on Siamese networks specifically for fish tracking, this approach has the potential to be used in the field.
Figure 1: Fish trajectory under different frame rates (This figure was adapted from [11]).
@@
Figure 2: Fish tracking method based on YoloV5 and SiamRPN++ (This figure was reproduced from [78]).
The emergence of transformer-based tracking methods has revolutionized object tracking, including fish tracking in aquaculture environments. Initially proposed for natural language processing tasks, transformers have been successfully adapted for computer vision tasks, including object detection and tracking [79]. Transformer-based trackers, such as TransTrack [80] and STARK [81], have demonstrated state-of-the-art performance on various tracking benchmarks. In the context of fish tracking, transformer-based methods have shown promising results. For instance, Li et al. [38] proposed a transformer-based multiple fish tracking model (TFMFT) to address the issue of instance loss in aquaculture ponds with complex background disturbances. This approach represents a shift from previous tracking-by-detection (TBD) methods [41] that relied on separate detection and embedding paradigms. Instead, the end-to-end Transformer architecture enables unified processing for both training and tracking. Building upon this work, Liu et al. [82] introduced FishTrack, an online multi-fish tracking model, that incorporates a specially designed encoder to encode historical position information of fish and automatically update spatiotemporal information in an auto-regressive manner. This approach allows for the fusion of spatiotemporal information of fish targets without manual feature selection, demonstrating robustness even under long-term occlusion conditions.
Despite these advancements, significant challenges remain across different tracking approaches. Identity-based methods like idTracker.ai [73] still face limitations in handling complex 3D environments and high-density scenarios. Detection-tracking frameworks struggle with environmental variations and often require retraining for different conditions, while Siamese networks, although promising, need further validation in real-world aquaculture settings. Transformer-based methods face computational complexity issues and require substantial training data, which can be scarce in aquaculture settings. They may also struggle with extended tracking sequences and lack inherent temporal modelling capabilities. Specific challenges include achieving accurate ID matching at high stocking densities and maintaining tracking stability during long-term monitoring. Furthermore, the potential for overfitting when the training data is limited or not diverse enough is a particular concern in aquaculture, where environmental conditions can vary widely. For example, TFMFT faces limitations in achieving accurate identity (ID) matching at high stocking densities and experiences some tracking losses during long-term monitoring.
Future research should focus on further adapting transformer architectures to address specific challenges, such as developing efficient training strategies to handle the limited availability of annotated fish-tracking datasets, improving performance in high-density scenarios, and enhancing long-term tracking stability. Additionally, efforts should be directed towards developing more efficient and lightweight models, improving their ability to handle long-term tracking and occlusions, and creating strategies to effectively train these models with limited data.